# ShaRP: A Novel Feature Importance Framework for Ranking

### v3

@@ 7.1. Datasets
name source # tuples # features
Tennis (ATP) (Joao Fonseca, 2024b) 86 6
CS Rankings (CSR) (Berger, 2023) 189 5
Times Higher Education (THE) (Joao Fonseca, 2024a) 1,397 5
Synthetic (SYN) this paper 2,000 2 or 3
Table 1. Datasets used in our experimental study.
@@
avg. time (sec)
dataset # tuples # features score rank pair
ATP 86 6 0.004 0.026 0.004
CSR 189 5 0.002 0.022 0.003
THE 1,397 5 0.011 0.423 0.007
SYN 2,000 3 0.002 0.126 0.003
Table 2. Running time of exact computation.
7.2. Efficiency and Approximation
(a) Speedup vs. sample size, THE
(b) Speedup vs. sample size, CSR
(c) Speedup vs. sample size, ATP
(d) Speedup vs. max coalition size, THE
(e) Speedup vs. maximum coalition size, CSR
(f) Speedup vs. maximum coalition size, ATP
7.2.1. Running time of exact computation
By definition, when computing feature importance, Shapley values use all possible coalitions of features and all items in the dataset. We refer to this as the exact computation of local feature-based explanations. In our first experiment, we measure the running time of the exact computation for rank, score, and pairwise QoI computation for 3 of the real and 1 of the synthetic datasets in Table 1. We include only 1 synthetic dataset because all synthetic datasets have the same number of points ($m=2,000$) and at most 3 features, and differ in the structure of the correlations within the data. However, these correlations do not impact the running time of exact computation. Finally, we do not include the running time for the top-$k$ QoI because it is based closely on rank QoI in our implementation, and the running times of generating explanations for these two QoIs are indistinguishable.
Further, we observe that the running time of rank and score QoIs increases both with the number of items ($m$ in Algorithm 1) and with the number of features $d$. This is as expected, since exact feature importance according to score and rank QoIs depends linearly on dataset size $\mathcal{D}$ (since $m=|\mathcal{D}|$ samples are taken for exact computation) and exponentially on the number of features $d$ (there are $12^{d}-$ 1 feature coalitions).
The computation of pairwise QoI only uses two points in the dataset (the points making up the pair), and so we would expect its running time to be insensitive to $m$ but exponential in $d$. Note, however, that both rank and pairwise QoI require computing the new rank of an item, or pair of items, upon intervention as part of Shapley value computation (line 7 in Algorithm 2 for rank QoI, and line 6 in Algorithm 3 for pairwise QoI). This rank computation can be implemented in a number of ways. In our implementation, this operation is linear in dataset size $m$, which explains, for example, that the running time of pairwise QoI for THE ($m=1,397$ and $d=5$) is higher than the running time for ATP ($m=86$ and $d=6$).
(a) Fidelity vs. maximum sample size, THE
(b) Fidelity vs. maximum sample size, CSR
(c) Fidelity vs. maximum sample size, ATP
(d) Fidelity vs. max coalition size, THE
(e) Fidelity vs. maximum coalition size, CSR
(f) Fidelity vs. maximum coalition size, ATP
7.2.2. Running time and quality of approximation
As discussed in Section 5.1, we implement two methods for approximating feature importance to reduce the running time: reduce the number of samples ($m$ in Algorithm 1) and restrict maximum coalition size ($|\mathcal{S}|$ in Algorithm 1). In addition to measuring the running time, we also measure and report explanation fidelity (Eq. 6 in Section 6.1) and method agreement (Eq. 9) to quantify the quality of the approximation. Figures 56,  7, and 8 present these results for the THE dataset, where exact computation was slowest and so there is most substantial room for improvement, as well as for the other two real datasets, CSR and ATP.
Figure 5a shows the speed-up when fewer samples $m$ are used for the computation. We observe that reducing the number of samples from 1378 (exact) down to 20, while keeping the maximum number of coalitions at 4 (largest possible for 5 features), speeds up rank QoI by a factor of 32, bringing the running time from 0.423 sec down to 0.013 sec. We see similar results for CSR 5b and ATP 5c, where speed-up is smaller because of the total number of features, but it is still significant, 6 for CSR (for 20 points down from 189) and 3 for ATP (for 20 points down from 86). Importantly, this performance gain does not come with a noticeable reduction in fidelity, which stays over 0.98 (out of 1) for all sample sizes in all experiments, Figure 6a6c.
Figures 5d5f and 6d6f show the speed-up and fidelity, respectively, when coalition size is bounded. We observe that speed-up is most substantial for all QoIs for coalitions of size 1, but that fidelity is lower for this approximation method. We observe fidelity values of at least 0.83 for rank, 0.91 for score, and 0.98 for pairwise. Fidelity further improves for rank and score for coalitions of size at most 2, reaching 0.87, 0.94, and 1, respectively. The difference between the approximation methods is expected because they affect different aspects of the sampling process. Sample size affects the number of samples we draw while maximum coalition size affects the number of features we can sample concurrently.
In Figures 7 and 8, we present the method agreement between the approximation and the exact computation for CSRankings (CSR). We omit method agreement results for the other datasets, where ShaRP  performs similarly. In 7a and 7b, we present agreement of the approximation when we vary sample size for the rank and the score QoI. We evaluate the agreement using the Jaccard Index (considering the top-2 features), Kendall’s tau distance, and the Euclidean distance of the feature vectors (converted to unit vectors). Here, we see that performance is similar for both QoIs. The Jaccard index is over 0.9 for any sample size, and is the distance metric with the worst performance for both QoIs. This is worth noting as shorter explanations are often considered more interpretable (Molnar, 2020). Agreement is similar or higher for all QoIs when we vary maximum coalition size, see Figure 8a8c.
In summary, reducing the number of samples and bounding coalition size are both effective at improving the running time, while retaining high explanation fidelity and method agreement in practice. Nonetheless, unless otherwise stated, we use exact computation in the remainder of this section to reduce variability in the reported results.
7.3. Qualitative Analysis
7.3.1. CS Ranking
In our next set of experiments, we generate and qualitatively analyze explanations for CS Rankings.
In Figure 9, we show feature contributions to rank and score QoI for the CS Rankings dataset, aggregated over 10% strata of the ranking. According to rank QoI in Figure 9a Systems is the most important feature in all strata, followed closely by AI. Both features have the most positive contributions in the top strata and the most negative in the bottom strata. Further, we observe that feature contributions to score are less informative than to rank: both capture the same relative feature importance for the top 20%; however, feature contributions become small and very similar as more items are tied or nearly-tied for their scores, making it nearly impossible to quantify feature importance across strata.
Figure 3, previewed in the introduction, presents the top-$k$ QoI for this dataset. The feature importance it shows is consistent with Figure 9b (score QoI), but the high positive impact of Systems on a department’s presence at the top-$k$ is even stronger pronounced. Additionally, unlike the score QoI, this top-$k$ QoI shows that Theory is impactful for getting into the top-$k$. Further, both top-$k$ and rank QoIs show that Systems and AI swap places in terms of importance, while score QoI does not detect this.
In Figure 3b3d we continue our analysis of the top-$k$ and consider the relative ranking of three universities: Georgia Tech in rank 5, Stanford in rank 6, and UMich in rank 7. We wish to understand why Georgia Tech is ranked higher than Stanford (Figure 3c), and why Stanford is ranked higher than UMich (Figure 3d). In both cases, Georgia Tech and UMich have lower values for all features except Systems. The Systems value of Georgia Tech is high enough to overcome the contributions of other features and rank it higher than Stanford. However, for UMich, we see that, while Systems is the most important feature in the top-10% stratum, it is not important enough to move UMich above Stanford. Note that adding the contributions for each feature adds up to the total difference in rank between each pair.
7.3.2. Score-based Ranking with Synthetic Data
@@
7.4. Comparison to Other Methods
In this section, we compare explainability methods using the metrics described in Section 6. We restrict our analysis to a subset of the methods described in Section 2. The focus of ShaRP and the strength of Shapley values is individual explanations, for this reason, we do not compare to the methods that focus on providing global explanations. This excludes Yang et al., and Gale and Marian (Yang et al., 2018; Gale and Marian, 2020). To compare to HIL (Yuan and Dasgupta, 2023), we adapt their code to work with any score-based ranker and real data. We do not compare to PrefShap (Hu et al., 2022) which works only for pairwise datasets that use a generalized preferential kernel function. We compare to HRE (Anahideh and Mohabbati-Kalejahi, 2022) but use only four of their internal methods as provided by their public code base. Finally, we compare to SHAP (Lundberg and Lee, 2017) and LIME (Ribeiro et al., 2016b), due to their wide use and availability, even though they are not designed for ranking.
Implementation of HIL
HIL  (Yuan and Dasgupta, 2023) is the only other method that recommends the usage of ranks as a profit function for individual explanations in ranking. While this method is not general, we are interested in comparing it with our rank QoI. This was not straightforward because the method is available as a web app that works only for linear weight scoring functions and datasets of two Gaussian features. To compare the rank-relevance contributions introduced in that paper to the rank QoI, we adapted their method using their definitions and code. This implementation is available alongside our own. Further, we extended their method to work with the specific non-linear scoring function used by CS Ranking, by changing the way that Std rank and Std score (discussed below) are computed.
@@ 7.4. Comparison to Other Methods
Details of HRE
For the purposes of comparison, we describe HRE (Anahideh and Mohabbati-Kalejahi, 2022) in more detail here. This method explains the ranked outcome of an item by fitting interpretable models in the immediate neighborhood of the item in the ranking (i.e., items that are ranked 5-10 consecutive positions above and below the item being explained). The interpretable models they make available through their public implementation are Decision Trees (DT), Linear Regression (LR), Ordinary Least Squares (OLS), and Partial Least Squares (PLS). Note that the feature-based explanations provided by HRE cannot be used to reconstruct the ranked outcome, and so we cannot compute fidelity for this method.
7.4.1. Fidelity
Table 3 presents the fidelity of all methods for which explanations can be used to reconstruct the ranked outcome, averaged across all items in all datasets. All methods are executed using their recommended settings to compute explanations for score QoI. Additionally, we include fidelity for ShaRP  for the rank QoI. Recall that ShaRP  is the only method that can compute an explanation for this QoI. We observe that LIME, SHAP and ShaRP  are all achieving high explanation fidelity, on average. HIL has reasonable fidelity for CSR but does not perform consistently on other datasets.
@@ 7.4.1. Fidelity
7.4.2. Agreement Between Explanations
Figure 13 presents the agreement between ShaRP  and all other methods for the rank and score QoIs in Figure 13a and Figure 13b, and the agreement between all methods averaged across all datasets in Figure 13c. We chose the rank and score QoIs for this comparison because we want to compare across all items of the dataset. and the methods we compare ShaRP to calculate neither pairwise nor presence in the top-k. Additionally, the results presented here are for Kendall’s tau distance between the explanations.
From these figures, we see that the explanations we will produce highly depend on the explanation method, regardless of the QoI. We see that ShaRP agrees the most with LIME and SHAP across both rank and score QoIs. HRE is expected to differ perhaps as it uses only localized information to produce explanations. However, when we are examining the similarity of explanations between HRE variants, we can still see that explanations are different for different HRE variants.
@@ 7.4.2. Agreement Between Explanations
(c) Method Agreement between all methods averaged across all datasets
@@
7.4.3. Sensitivity
@@
Figure 14. Sensitivity analysis for the CSRankings dataset for all methods. Each dot is a neighbor of the reference item. The x-axis shows the Euclidean explanation distance between the item and each neighbor, the y-axis the absolute rank difference between the reference item and each neighbor, and the hue shows how similar the features of each neighbor are to the reference item. ShaRP is the only method that can successfully provide similar explanations for items that are similar in feature space, and that also have similar outcomes.
Figure 14 presents the results of sensitivity analysis for all methods. In this figure, we evaluate the similarity of explanations for pairs of similar items. For each pair of items, we compute three distances: (1) Euclidean distance between the explanations (x-axis); (2) distance between the ranks of the two items (y-axis); and (3) Euclidean distance between the items in terms of their feature values (hue, where lighter means closer). To make the plot, we place one item (the reference item) at position (0,0) and use a scatter point for each other item (neighbor), indicating the distance in ranks and the distance of the explanations. The color of the scatter point indicates the distance between the features of the reference point and the neighbor. We then overlay the plots for all items in the dataset, so that all items are used as reference points.
Intuitively, we would expect items that are both similar in terms of their features and ranked near each other to have similar explanations. In other words, we would expect all points to be on or near the diagonal line $y=x$, with the hue getting darker as we move away from the origin. However, this is rarely the case in practice. For example, if one feature is far more important than the rest, then proximity in the Euclidean space of features will not be strongly correlated with explanation similarity.
Another example where this intuition does not hold is when two items that look very different in the feature space receive very similar outcomes. In both cases, we would like the explanations to be able to show this. Indeed, we should desire different explanations for items ranked close by, when their features are different. And we never want any explanations of items that are far from each other in the ranking to be very similar. This means that we want the plot to fill the space under the $y=x$ line and have a hue that gets darker as we move further from the origin.
Consider Figure 14, and observe that both ShaRP  with rank QoI (Figure 14h) and HIL-Std-Rank (Figure 14j, using our expanded definition of their method) are indeed able to produce similar explanations for similarly ranked items that have similar features, conforming to the shape we described above.  ShaRP  is performing slightly better in clustering the points into denser similarity groups. On the other hand, SHAP (Figure 14f), a score-based method, bases its explanations mostly on the feature distances. Indeed, the plot consists of clear bands that get darker away from 0 and below the diagonal. Nevertheless, items that are ranked very differently but have similar features receive almost identical explanations. ShaRP  with score QoI in Figure 14g is almost identical to SHAP, as we also saw in method agreement, and HIL using the expanded score definition in Figure 14i also performs similarly. LIME (Figure 14e), which is also score-based, is producing a plot that suggests that the method uses both feature distance and ranking distance, but is still unable to differentiate between the explanations as clearly as the rank-based methods. Finally, the explanation distance of all HRE methods does not appear to depend on the ranking distance or the feature distance as we see items with similar and dissimilar explanations across all ranks and hues. This is expected, as the explanations depend only on close neighbors (that can have different features) and there is no guarantee that a different cluster of explanations for a further neighborhood will be different from the neighborhood of the reference point.
To further compare ShaRP  with rank QoI and HIL with Std rank, we present Figure 15. Even though both methods are appropriate for the ranking task we are examining, in this figure, we see that ShaRP  with rank QoI Figure 15a,  15c, and  15e can capture the full range of different ranks and features, and that groups the items more successfully. HIL with Std rank cannot capture the difference of feature values for ATP (Figure 15b), or the similarly ranked items that have different feature values for the Synthetic experiment (the middle area close to the $x$-axis of Figure 15f. Both methods perform similarly for THE (Figures 15c and  15d).
@@
8. Conclusions, Limitations, and Future Work
We presented ShaRP , a comprehensive framework for quantifying feature importance for selection and ranking. Because of the impact rankers have on individuals, organizations, and population groups, there is a need to understand them: to know whether the decisions are abiding by the law (auditing and compliance tasks), to help individuals improve their ranked outcomes (recourse tasks), and to design better ranking procedures (design task). Our work is motivated by the foundational interpretability needs of these tasks.
We showcased the usefulness of ShaRP  for qualitative analysis of an impactful real dataset and task — the ranking of Computer Science departments. We augmented this analysis with a thorough evaluation of our methods on numerous real and synthetic datasets, showing that the quantities of interest we define bring interesting — and complementary — insights about the data that cannot be captured simply by the relationship between the scores and the ranks. We showed that feature importance depends on the properties of the data distribution even if the scoring function remains fixed, and, further, that feature importance exhibits locality. Finally, we compared our method to other local feature-based explanation methods and showed that it compares to them favorably.
Limitations and future work.
A limitation of our work is that our evaluation of the usefulness of explainability did not include a user study. Understanding how humans make sense of feature importance in ranking and act on their understanding, is in our immediate plans.
We also plan to support additional quantities of interest, incorporate learning-to-rank methods into our evaluation, and develop a comprehensive benchmark that compares both local and global feature importance methods for ranking in terms of their usability, expressiveness, and performance.
Acknowledgements.
We would like to thank Tilun Wang for her contributions to the code base that supported an earlier version of this draft. We would also like to thank Lucius Bynum and Falaah Arif Khan for their input and advice while developing this work. This research was supported in part by NSF Awards No. 2312930 and 2326193, and by the NYU Tandon Center for Responsible AI, through their Responsible AI for Ukraine program, r-ai.co/ukraine.

### v4

@@ A.1. Datasets
(e) Moving Company - train
(f) Moving Company - test
@@
Moving Company
The moving company scenario (Yang et al., 2021) simulates a hiring process where job applicants are ranked based on their qualification score, computed as a function of their weight lifting ability, sex, and race. We train two different rankers, over two scenarios:
  1. (1)
    Using the original data from a previous hiring process from that company, where female applicants generally display lower weight-lifting ability than male applicants and a lower qualification score. In addition, black applicants have a lower qualification score compared to white applicants, but similar weight-lifting ability. Hence black females face greater discrimination compared to the rest of the applicants.
  2. (2)
    After applying the intersectional fairness intervention proposed in the same paper over the data.
All versions of this dataset (both scenarios and train/test sets) contain 2000 tuples.
We use an eXtreme Gradient Boosting (XGB) and a Light Gradient-Boosting (LGB) Machine to model the rankings of the applicants in the training set, and infer and compute the feature contributions of the applicants in the test set, i.e., all results reported in this section correspond to the test set. The XGB ranker was defined with the pairwise ranking objective, while the LGB ranker uses the lambdarank objective.
@@
Appendix B Distributional Analysis for Ranking
@@
Appendix C Implementation of HIL
HIL  (Yuan and Dasgupta, 2023) is the only other method that recommends the usage of ranks as a profit function for individual explanations in tabular ranking. While this method is not general, we are interested in comparing it with our rank QoI. This was not straightforward because the method is available as a web app that works only for linear weight scoring functions and datasets of two Gaussian features. To compare the rank-relevance contributions introduced in that paper to the rank QoI, we adapted their method using their definitions and code. This implementation is available alongside our own. Further, we extended their method to work with the specific non-linear scoring function used by CS Ranking, by changing the way that Std rank and Std score (discussed below) are computed.
@@ Appendix C Implementation of HIL
Appendix D Additional Details on Method Comparisons
D.1. Fidelity
We provide more details on the Fidelity results discussed in Section 8.2.2. We compute the Fidelity of all the methods that have that property across all datasets. We use SHAP and LIME out-of-the-box so their performance is not perfect (although extremely good). We make this choice to highlight the importance of using exact Shapley values when computing local explanations, where the error in each separate explanation is important as each explanation impacts a separate person.
@@ D.1. Fidelity
D.2. Agreement between Explanations
Figure 13 provides a per-dataset visualization of the agreement between the explanations of the methods in Section 8.2.3.
In Fig. 13a We visualize Kendall’s tau explanation distance correlation of ShaRP using the rank QoI with all other methods across every dataset. In Fig. 13b, we plot the same result for ShaRP using the score QoI. As in the aggregated method agreement plot (Fig. 7), ShaRP is very similar to SHAP and LIME for both QoIs. As expected it is more similar to the score QoI but not identical that is perhaps because we used SHAP out-of-the-box which applies some approximation parameters for running time optimization. Similarly, ShaRP behaves similarly to what we discussed in Fig. 7 to all methods across the datasets except the HIL methods for the Synthetic datasets. We hypothesize that this is because the HIL methods are able to perform better for those datasets due to their Synthetic nature.
@@ D.2. Agreement between Explanations
D.3. Sensitivity
We provide the sensitivity analysis for all methods for CSRankings in Fig. 14. In addition, to HRE LR, DEXER, LIME, SHAP, ShaRP rank, and HIL std rank that we presented in Fig. 6, we also plot HRE DT in Fig. 14a, HRE OLS in Fig. 14c, HRE PLS in Fig. 14d, ShaRP score in Fig. 14g, HIL std score in Fig. 14i. We see that all HRE methods perform similarly or worse than HRE-LR, this is unsurprising as all these methods are used locally. We also see that both HIL std score and ShaRP score perform similarly to SHAP, this is also expected. HIL std score and DEXER are very similar which reveals our inability to predict the rank using the ranked output of the model. Specifically, HIL std score assumes knowledge of the weights used by the model and uses them directly to compute the feature importance. DEXER is assuming black-box access to the ranked output only and fits a linear regression model on the ranking. Nevertheless, judging from these results, it appears that DEXER is explaining the score (and not the rank) and is learning the model weights to do so. Inadvertently, we also show that the choice of the explanation method makes a big difference to the final explanation.
@@ D.3. Sensitivity
(k) DEXER
Figure 14. Sensitivity analysis for the CSRankings dataset for all methods
To further compare ShaRP  with rank QoI and HIL with Std rank, we present Figure 15. Even though both methods are appropriate for the ranking task we are examining, in this figure, we see that ShaRP  with rank QoI (Figures 15a,  15c, and  15e) can capture the full range of different ranks and features, and that groups the items more successfully. HIL with Std rank cannot capture the difference of feature values for ATP (Figure 15b), or the similarly ranked items that have different feature values for the Synthetic experiment (the middle area close to the $x$-axis of Figure 15f). Both methods perform similarly for THE (Figures 15c and  15d).
@@
Finally, we present an analysis of ShaRP using the score QoI and the rank QoI for the CSR dataset but for a score task (instead of rank). The goal of this analysis is to show that the sensitivity of the methods that use a score QoI is very high when we are explaining a score task. In other words, if we are trying to explain the score, then the methods that use a score-based profit function perform the best as it is fully expected.
The task we are going to explain is the score of the CSRankings scoring function. We choose this task for two reasons, first, we already provided the results of the CSRankings ranking task and we can draw a direct comparison. Secondly, we have a ranking for that dataset and we can plot the methods that use the rank QoI for juxtaposition. Note that it is entirely redundant to use a rank-based QoI method in this case. In fact, it is redundant to even produce a ranking as we are asking an explainability question about the score. But we are choosing to provide this information to showcase that each explainability task needs its own profit function and the choice of the profit function makes a big difference to the final explanation.
In Fig. 16, we evaluate the similarity of explanations for pairs of similar items when we attempt to explain the score. For each pair of items, we compute three distances: (1) Euclidean distance between the explanations (x-axis); (2) distance between the scores (instead of rank) of the two items (y-axis); and (3) Euclidean distance between the items in terms of their feature values (hue, where lighter means closer). To make the plot, we place one item (the reference item) at position (0,0) and use a scatter point for each other item (neighbor), indicating the distance in ranks and the distance of the explanations. The color of the scatter point indicates the distance between the features of the reference point and the neighbor. We then overlay the plots for all items in the dataset, so that all items are used as reference points.
Unlike Fig. 14, we now expect to see items that are both similar in terms of their features and scored near each other to have similar explanations. We would still expect all points to be on or near the diagonal line $y=x$, with the hue getting darker as we move away from the origin, if their explanations successfully explain the score.
In Fig. 16, we see that indeed the score-based methods have the desired shape we discussed in Section 8.2.1. ShaRP score is extremely similar and almost entirely fits the $y=x$ line. ShaRP rank, appears to be providing explanations that do not depend on the score distance between the items’ outcomes (y-axis) or the feature distance between the items (hue) as expected.
This analysis shows how QoI selection is important when providing an explanation. The score is unable to perform well for a ranking task since it estimates the impact of each feature on the score outcome and similarly, it is completely unreasonable to use a rank QoI when explaining the score.
(a) ShaRP Score
(b) ShaRP Rank
Figure 16. Sensitivity analysis for the CSRankings dataset for all methods when the task we are trying to explain is a score task. Compared to Figure 14 we see that the methods that are using a score QoI are performing better.
(a) Speedup vs. sample size, CSR
(b) Speedup vs. sample size, ATP
(c) Speedup vs. maximum coalition size, CSR
(d) Speedup vs. maximum coalition size, ATP
Figure 17. Computational time performance of approximation for CSRankings (CSR), and ATP Tennis (ATP). Speedup is computed in comparison to exact computation times, reported in Table 2.
(a) Fidelity vs. maximum sample size, THE
(b) Fidelity vs. maximum sample size, CSR
(c) Fidelity vs. maximum sample size, ATP
(d) Fidelity vs. maximum coalition size, CSR
(e) Fidelity vs. maximum coalition size, ATP
D.4. Additional Efficiency and Approximation Additional Results
In this section, we present the extended results previewed in Section 8.3.2.
In Figure 17 we present the speedup vs. sample size, and speed-up vs max coalition size for THE, CSR, and ATP. We already presented the results for THE in Figure 8. We observe similar results but scaled down due to the dataset sizes. In Figure 18, we present the corresponding fidelity for both sample size and max coalition size. In this plot we also provide the fidelity of THE for sample size that we omitted in the main body of this work due to space constraints. We observe that fidelity is very high for all sample sizes, and almost identical to the fidelity of THE for all max coalition sizes.
In Figures 19 and 20, we present the method agreement between the approximation and the exact computation for CSRankings (CSR). We omit method agreement results for the other datasets, where ShaRP  performs similarly. In 19a and 19b, we present agreement of the approximation when we vary sample size for the rank and the score QoI. We evaluate the agreement using the Jaccard Index (considering the top-2 features), Kendall’s tau distance, and the Euclidean distance of the feature vectors (converted to unit vectors). Here, we see that performance is similar for both QoIs. The Jaccard index is over 0.9 for any sample size, and is the distance metric with the worst performance for both QoIs. This is worth noting as shorter explanations are often considered more interpretable (Molnar, 2020). Agreement is similar or higher for all QoIs when we vary maximum coalition size, see Figure 20a20c.